# AutoGuide: Automated Generation and Selection of State-Aware Guidelines for Large Language Model Agents

### v1

@@ Abstract
The primary limitation of large language models (LLMs) is their restricted understanding of the world. This poses significant difficulties for LLM-based agents, particularly in domains where pre-trained LLMs lack sufficient knowledge. In this paper, we introduce a novel framework, called AutoGuide, that bridges the knowledge gap in pre-trained LLMs by leveraging implicit knowledge in offline experiences. Specifically, AutoGuide effectively extracts knowledge embedded in offline data by extracting a set of state-aware guidelines. Importantly, each state-aware guideline is expressed in concise natural language and follows a conditional structure, clearly describing the state where it is applicable. As such, the resulting guidelines enable a principled way to provide helpful knowledge pertinent to an agent’s current decision-making process. We show that our approach outperforms competitive LLM-based baselines by a large margin in sequential decision-making benchmarks.
[…]
[0]leftmargin=*
@@ 1 Introduction
Recent advances in large language models (LLMs) have empowered AI agents to address various sequential decision-making tasks and applications (Wang et al., 2023; Xi et al., 2023). The foundation of these successes involves the planning and reasoning capabilities of pre-trained LLMs, enabling agents to execute effective policies (Brohan et al., 2023; Wei et al., 2022). However, the knowledge embedded in pre-trained LLMs is confined to the concepts and information present in a training dataset, which is inherently incomplete (Yin et al., 2023). This limited understanding in pre-trained LLMs poses significant difficulties for agents, particularly in target domains with insufficient prior knowledge. For example, in web navigation, LLM agents generally achieve low success rates due to a lack of information about taking appropriate actions on real-world websites with diverse and dynamic contents (Koh et al., 2024; Deng et al., 2023; Gur et al., 2023; Zhou et al., 2023).
@@
In this paper, we explore data-driven strategies that leverage offline experiences to bridge knowledge gaps between pre-trained LLMs and a downstream domain. As offline experiences implicitly convey valuable knowledge about desirable and undesirable policies in domains, they promise to serve as a useful resource for improving an LLM agent’s decision-making in situations where the pre-trained LLM lacks understanding. Despite this potential benefit, a critical challenge lies in effectively extracting the implicit information embedded in offline data. Unfortunately, the conventional method of directly providing all available experiences to an agent, as in in-context learning, is either impossible or unsuccessful due to constraints, including context length limitations, prompt sensitivity, and difficulty with complex reasoning (Lu et al., 2022; Dong et al., 2022; Min et al., 2022; Kaddour et al., 2023).
@@
To address the challenge of extracting knowledge from offline data, we propose a novel framework, called AutoGuide. Specifically, AutoGuide automatically extracts a comprehensive set of state-aware guidelines from offline experiences. Our method then applies these state-conditional guidelines to improve an LLM agent’s performance by identifying the current state and then incorporating its corresponding guidelines into the prompt during testing (see Figure 1). Notably, we generate state-aware guidelines in concise natural language statements, efficiently compressing knowledge in offline data. Additionally, state-aware guidelines clearly describe the states where each guideline is applicable, so AutoGuide provides an LLM agent with guidelines that are most relevant to its current decision-making process. As a result, AutoGuide achieves the highest success rates compared to competitive baselines in challenging sequential decision-making benchmark environments.
@@
Figure 1: AutoGuide aims to extract the implicit knowledge embedded in offline experiences and help the decision-making process of an LLM agent. Specifically, our method generates a comprehensive set of state-aware guidelines from offline data and explicitly identifies when each guideline is applicable by generating its corresponding state. Our state-aware guidelines enable providing pertinent guidelines at test time by summarizing the current trajectory into the state, leading to correct decision-making compared to baselines without state-aware guidelines.
@@
  • Principled method based on state-aware guidelines (Section 3): We develop two modules to generate state-aware guidelines from offline experiences: the state summarization module for generating a description of a given state, and the guideline extraction module for extracting a desired guideline corresponding to the state. The outcome is a set of domain knowledge in concise natural language that enhances decision making by providing pertinent information.
  • Comprehensive evaluation of AutoGuide (Section 4.2): We show AutoGuide’s capability in extracting helpful state-aware guidelines in various interactive benchmark domains, including navigating realistic web environments. Our results highlight the flexibility of AutoGuide, such that it can be seamlessly incorporated into an existing LLM-based method.
  • Analyses with important perspectives (Section 4.3): We explore various aspects of AutoGuide, such as the significance of determining the applicability of each guideline based on generated states. We also investigate having only successful or unsuccessful trajectories in offline experiences, and demonstrate that AutoGuide can still extract valuable knowledge in these corner cases.
@@ 2 Related Work
LLM agents. Language models have recently been shown to possess strong priors for sequential decision-making tasks, which has given rise to LLM-powered agents (Wang et al., 2023; Xi et al., 2023; Zheng et al., 2024; Zeng et al., 2023). Agents need to possess various skills to be effective in practice including planning (Brohan et al., 2023; Huang et al., 2022; Logeswaran et al., 2022), reasoning (Wei et al., 2022; Gao et al., 2023), tool manipulation (Qin et al., 2023; Patil et al., 2023; Parisi et al., 2022; Schick et al., 2023), code generation (Sun et al., 2023; Logeswaran et al., 2022), among others. In this work we focus on building effective agents for web (Yao et al., 2022b; Zhou et al., 2023) and embodied (Shridhar et al., 2021) environments.
@@
Self-reflection from past experiences. An important capability for agents to succeed is the ability to learn from past experiences and update their behavior based on feedback. Self-feedback (Madaan et al., 2023; Kim et al., 2023; Shinn et al., 2023) has emerged as an effective technique where a model inspects its own incorrect predictions, reflects on it to identify what went wrong and attempts to improve its prediction. Our approach shares similar motivations to self-refinement approaches, but unlike these methods which generate reflections on the fly, our approach first attempts to distill a set of guidelines from offline experiences and then retrieves the appropriate guidelines during inference for more accurate action prediction. However, these advances are complementary to our method and can be used in conjunction with our approach (as shown in the experiments).
@@
Leveraging natural language guidance. Natural Language can be a rich source of information for artificial agents to learn to act efficiently. Prior work has explored the notion of learning from human written text manuals which describe various details about the agent’s environment (Branavan et al., 2012; Hanjie et al., 2021; Zhong et al., 2020). Recent work has explored automatically generating such guidance in the form of chain-of-thought reasoning (Wei et al., 2022; Yao et al., 2022b) which emulates a thought process or rationale for an agent’s predictions. In contrast to ReAct, which generates such guidance on the fly during inference, our approach examines successful and unsuccessful trajectories corresponding to a task to generate appropriate guidance and uses these guidelines for predicting better actions. ExpeL (Zhao et al., 2023) proposed a related approach to derive guidelines. In contrast to ExpeL, where all guidelines are provided to an agent as a prompt, our guideline selection process is contextual, where guidelines relevant to the agent’s current state are retrieved and used for prediction. We show that this substantially improves over ExpeL’s non-contextual guideline-based approach.
Figure 2: State-aware guideline generation process based on a pair of success trajectory $\bm{\tau}^{i}_{\text{success}}$ and failure trajectory $\bm{\tau}^{i}_{\text{fail}}$. In this example, the two trajectories start deviating from each other at $t=1$. The state summarization module generates a description of a state at $t=1$ given $\bm{\tau}^{i}_{:t}$, and the guideline extraction module generates the corresponding guideline for that state.
3 AutoGuide: Principled Method Based on State-Aware Guidelines
@@
Our work is motivated by the growing availability of offline experiences that agents or humans naturally accumulate during their interactions with the environment. We focus on leveraging these offline experiences to enhance the decision-making of an LLM agent. A key question in achieving our goal is how to effectively extract implicit knowledge embedded in these experiences. This section details our method, AutoGuide, which extracts a comprehensive set of state-aware guidelines from offline data.
@@ 3.1 Problem Statement
The goal of AutoGuide is to distill knowledge from offline data into a helpful natural language format, such that the extracted information improves the decision-making of an agent during testing. Formally, AutoGuide is given offline experiences $\mathcal{D}_{\text{train}}\!=\!(\bm{\tau}^{1},...,\bm{\tau}^{N})$ that consist of $N$ trajectories from the training tasks. Each trajectory $\bm{\tau}^{i}\!:=\!(x_{0},a_{0},r_{0},...,r_{T})$ is a sequence of observations, actions, and rewards following the Markov decision process (Sutton & Barto, 2018). We assume that each trajectory $\bm{\tau}^{i}$ can be categorized as either a success or failure determined by either direct feedback from the environment or the terminal reward (e.g., considered a success if the final reward $r_{T}$ is positive) depending on the domain.
3.2 Constructing Dataset for Guideline Extraction
From an offline experience dataset $\mathcal{D}_{\text{train}}$, we construct a paired offline dataset $\mathcal{D}_{\text{paired}}$ for guideline extraction:
$\begin{split}\mathcal{D}_{\text{paired}}\!&=\!\big{(}(\bm{\tau}^{1}_{\text{success}},\bm{\tau}^{1}_{\text{fail}}),...,(\bm{\tau}^{n}_{\text{success}},\bm{\tau}^{n}_{\text{fail}})\big{)},\end{split}$ (1)
where $\bm{\tau}^{i}_{\text{success}}\!\in\!\mathcal{D}_{\text{train}}$ and $\bm{\tau}^{i}_{\text{fail}}\!\in\!\mathcal{D}_{\text{train}}$ denote a success and failure trajectory in $\mathcal{D}_{\text{train}}$ of the same task, respectively. In practice, we can easily collect these pairs by applying advanced LLM-based planning approaches (Yao et al., 2022b; Shinn et al., 2023) or taking advantage of human demonstration datasets (Zhou et al., 2023; Deng et al., 2023). However, in rare cases, it can be challenging to obtain paired data if the offline data $\mathcal{D}_{\text{train}}$ contains only one type (either success or failure) of the trajectories. We discuss how our algorithm can be adjusted to handle such a biased dataset and its impact on the performance in Section 4.3.
3.3 Extraction of State-Aware Guidelines
@@
AutoGuide generates state-aware guidelines by utilizing pairs of success and failure trajectories from offline data. Each state-aware guideline is expressed in concise natural language and follows a conditional structure, clearly describing the state in which this guideline is applicable. Intuitively, contrasting actions taken between $\bm{\tau}^{i}_{\text{success}}$ and $\bm{\tau}^{i}_{\text{fail}}$ provides important information about when and which actions are effective or ineffective. Building on this insight, we develop two modules for deriving state-aware guidelines in a conditional structure (see Figure 2):
Algorithm 1 Collection of state-aware guidelines
@@
  Input: Paired offline dataset $\mathcal{D}_{\text{pair}}$, state summarization module $\mathcal{M}_{\mathrm{state}}$, guideline extraction module $\mathcal{M}_{\mathrm{guideline}}$
  Initialize state-aware guideline dictionary $\mathcal{G}$
  for Each pair ($\bm{\tau}^{i}_{\text{success}},\bm{\tau}^{i}_{\text{fail}})\in\mathcal{D}_{\text{pair}}$  do
     # Summarize a trajectory into a state
     Identify target timestep $t$ from $\bm{\tau}^{i}_{\text{success}}$ and $\bm{\tau}^{i}_{\text{fail}}$
     $\texttt{state}\leftarrow\mathcal{M}_{\mathrm{state}}(\bm{\tau}^{i}_{:t})$
     # Check if the current state matches any existing states
     if $\texttt{state}\notin\mathcal{G}$ then
        $\mathcal{G}[\texttt{state}]=\{\}$
@@
     # Generate the corresponding state-aware guideline
     $\texttt{guideline}\leftarrow\mathcal{M}_{\text{guideline}}(\bm{\tau}^{i}_{\text{success}},\bm{\tau}^{i}_{\text{fail}},\texttt{state})$
     $\mathcal{G}[\texttt{state}]\leftarrow\mathcal{G}[\texttt{state}]\cup\{\texttt{guideline}\}$
@@
  Return State-aware guideline dictionary $\mathcal{G}$
State summarization module. The objective of this module is to generate a concise description of the state of a trajectory in natural language. Specifically, given a pair of $\bm{\tau}^{i}_{\text{success}}$ and $\bm{\tau}^{i}_{\text{fail}}$ from the same task, we first identify the target timestep $t$ at which these two trajectories start deviating from each other with different actions. Then, we construct a target trajectory $\bm{\tau}^{i}_{:t}\!:=\!(x_{0},a_{0},...,x_{t})$, which comprises a sequence up to $t$, either from a success or failure trajectory and prompt an LLM to summarize it:
$\begin{split}\text{{state}}\!\leftarrow\!\mathcal{M}_{\text{state}}(\bm{\tau}^{i}_{:t}),\end{split}$ (2)
where Section B.1 details our prompt template. For example, Figure 2 shows trajectories related to the task $i$ of creating a discussion post about online learning in a relevant subReddit. Given that these two trajectories have different actions at a target timestep of $t\!=\!1$, our state summarization module generates the following state summary from $\bm{\tau}^{i}_{:1}$ in Figure 2: “You are on the List of forums page”.
Guideline extraction module. The goal of this module is to generate a desired guideline corresponding to the state. We extract a useful natural language guideline by examining successful and failure trajectories with respect to the state:
$\begin{split}\text{{guideline}}\!\leftarrow\!\mathcal{M}_{\text{guideline}}(\bm{\tau}^{i}_{\text{success}},\bm{\tau}^{i}_{\text{fail}},\texttt{state}),\end{split}$ (3)
where we refer to Section B.2 for the prompt template we use. As an example, this module generates the following state-aware guideline for the state described in Figure 2: “When you are on the List of forums page, … you can click on the link that exactly matches the forum name …”.
Algorithm 2 Applying state-aware guidelines at test time
  Input: State-aware guideline dictionary $\mathcal{G}$, state summarization module $\mathcal{M}_{\text{state}}$, guideline selection module $\mathcal{M}_{\text{select}}$, LLM agent policy $\pi$
@@
     # Summarize a trajectory into a state
     $\texttt{state}\leftarrow\mathcal{M}_{\text{state}}(\bm{\tau})$
     # Check if the current state matches any existing states. If matched, then perform top-$k$ guideline selection
     if $\texttt{state}\in\mathcal{G}$ then
        $\texttt{guidelines}\leftarrow\mathcal{M}_{\text{select}}(\texttt{state},\mathcal{G}[\texttt{state}],\bm{\tau})$
@@
        $\texttt{guidelines}\leftarrow\varnothing$
@@
     $a_{t}\sim\pi(\bm{\tau},\texttt{state},\texttt{guidelines})$
@@
     Update trajectory $\bm{\tau}\leftarrow\bm{\tau}\cup\{\texttt{state},a_{t},x_{t+1}\}$
@@
Construction of state-aware guidelines. We collect state-aware guidelines $\mathcal{G}$ by iterating through available pairs in the paired offline data and organize the guidelines in a dictionary format, using the state as the key and the corresponding guidelines as the value (see Algorithm 1). In particular, we observe that the state summarization module occasionally produces states that share the same context but are expressed slightly differently. To minimize redundancy, we employ an LLM to determine if the current state corresponds to any previously identified state. If a match is found, we reuse the existing state; otherwise, we introduce the new state into our dictionary. The specific prompt template for this state-matching procedure is outlined in Section B.3.
Figure 3: Sequential decision-making benchmark domains considered in our work: ALFWorld (Shridhar et al., 2021), WebShop (Yao et al., 2022a), WebArena (Zhou et al., 2023). Graphic credit: Shridhar et al. (2020); Yao et al. (2022a); Zhou et al. (2023).
3.4 Applying State-Aware Guidelines at Test Time

### v2

# AutoGuide: Automated Generation and Selection of Context-Aware Guidelines for Large Language Model Agents

@@ Abstract
Recent advances in large language models (LLMs) have empowered AI agents capable of performing various sequential decision-making tasks. However, effectively guiding LLMs to perform well in unfamiliar domains like web navigation, where they lack sufficient knowledge, has proven to be difficult with the demonstration-based in-context learning paradigm. In this paper, we introduce a novel framework, called AutoGuide, which addresses this limitation by automatically generating context-aware guidelines from offline experiences. Importantly, each context-aware guideline is expressed in concise natural language and follows a conditional structure, clearly describing the context where it is applicable. As a result, our guidelines facilitate the provision of relevant knowledge for the agent’s current decision-making process, overcoming the limitations of the conventional demonstration-based learning paradigm. Our evaluation demonstrates that AutoGuide significantly outperforms competitive baselines in complex benchmark domains, including real-world web navigation.
@@ 1 Introduction
Recent advances in large language models (LLMs) have empowered AI agents to address various sequential decision-making tasks and applications [1, 2]. The foundation of these successes involves the planning and reasoning capabilities of pre-trained LLMs, enabling agents to execute effective policies [3, 4]. The predominant approach to leveraging these (typically closed source) models for sequential decision making tasks is to provide demonstrations in the form of in-context examples. However, direct application of this learning paradigm can be limited, especially in target domains where the LLM has insufficient prior knowledge such as in web navigation, where LLM agents generally achieve low success rates due to diverse and dynamic contents [5, 6, 7, 8]. Providing all available experiences as demonstrations to an agent can further be unsuccessful due to context length limitations, prompt sensitivity, and difficulty with complex reasoning [9, 10, 11, 12].
@@
On the other hand, LLMs excel in interpreting concise instructions provided as natural language, an ability that is also reinforced in the instruction-tuning phase of LLMs. Inpired by this, we explore data-driven strategies that leverage offline experiences to extract actionable knowledge to help guide LLM agents. As offline experiences implicitly convey valuable knowledge about desirable and undesirable policies in domains, they promise to serve as a useful resource for improving an LLM agent’s decision-making in situations where the pre-trained LLM lacks understanding. Despite this potential benefit, a critical challenge lies in effectively extracting the implicit information embedded in offline data.
@@
To address the challenge of extracting knowledge from offline data, we propose a novel framework, called AutoGuide. Specifically, AutoGuide automatically derives a comprehensive set of context-aware guidelines from offline experiences. Our method applies these context-conditional guidelines to enhance the performance of an LLM agent by retrieving guidelines relevant to the agent’s current state and incorporating them into the prompt during testing (see Figure 1). Notably, we generate context-aware guidelines in concise natural language statements, effectively compressing knowledge in offline data. Moreover, context-aware guidelines clearly describe the contexts where they are applicable, so AutoGuide enables an LLM agent to select pertinent guidelines for its current decision-making process. As a result, AutoGuide achieves the highest success rates compared to competitive baselines in complex sequential decision-making benchmark environments.
@@
Figure 1: AutoGuide aims to extract the implicit knowledge embedded in offline experiences and help the decision-making process of an LLM agent. Specifically, our method generates a comprehensive set of context-aware guidelines from offline data and explicitly identifies when each guideline is applicable by generating its corresponding context. Our context-aware guidelines enable providing pertinent guidelines at test time by identifying the context of the current trajectory, leading to correct decision-making compared to baselines without context-aware guidelines.
@@
  • Principled method based on context-aware guidelines (Section 3): We develop two modules to automatically generate context-aware guidelines from offline experiences: the context identification module for identifying the context of a given trajectory, and the guideline extraction module for extracting a desired guideline corresponding to that context. The outcome is a set of domain knowledge in concise natural language that enhances decision-making by providing pertinent information.
  • Comprehensive evaluation of AutoGuide (Section 4.2): We show AutoGuide’s capability in extracting helpful context-aware guidelines in various interactive benchmark domains, including navigating real-world web domains (e.g., GitHub). Our results highlight the effectiveness of AutoGuide, which significantly outperforms baselines without context-aware guidelines.
  • Analyses with important perspectives (Section 4.3): We study various aspects of AutoGuide, such as the significance of determining the applicability of each guideline based on generated contexts. We also investigate the generalization ability of context-aware guidelines and demonstrate that our guidelines enhance the performance across out-of-domain tasks.
@@ 2 Related Work
LLM-based agents. Language models have recently been shown to possess strong priors for sequential decision-making tasks, which has given rise to LLM-powered agents [1, 2, 13, 14]. Agents need to possess various skills to be effective in practice including planning [3, 15, 16], reasoning [4, 17], tool manipulation [18, 19, 20, 21], code generation [22, 16], among others. In this work, we focus on building effective agents for web [23, 8] and embodied [24] environments.
@@
Self-reflection from past experiences. An important capability for agents to succeed is the ability to learn from past experiences and update their behavior based on feedback. Self-feedback [25, 26, 27] has emerged as an effective technique where a model inspects its own incorrect predictions, reflects on it to identify what went wrong and attempts to improve its prediction. While self-feedback provides intra-task (i.e., per-episode) knowledge within a task based on immediate feedback, our approach offers an orthogonal and complementary aspect of inter-task knowledge (over multiple tasks) by considering multiple train tasks in offline data. AutoGuide enhances learning efficiency and credit assignment by utilizing detailed feedback from multiple tasks. However, these self-feedback approaches are complementary to AutoGuide and can be used in conjunction with our approach, as shown in our experiments (see Section 4.2).
@@
Leveraging natural language guidance. Natural Language can be a rich source of information for agents to learn to act efficiently. Prior work has explored the notion of learning from human-written text manuals, which describe details about the environment [28, 29, 30]. Recent work has explored automatically generating such guidance in the form of chain-of-thought reasoning [4, 23], which emulates a thought process or rationale for agent’s predictions. In contrast to approaches which generate such guidance dynamically on the fly by imitating example guidance demonstrations provided by a human, our approach carefully compares trajectories in offline data to generate appropriate guidance and uses these guidelines for predicting better actions. ExpeL [31] proposed a related approach to derive guidelines. In contrast to ExpeL, where all guidelines are provided to an agent as a prompt, our guideline selection process is contextual, where guidelines relevant to the agent’s current state are retrieved and used for prediction. We show that this substantially improves over ExpeL’s non-contextual guideline-based approach.
3 AutoGuide: Principled Method Based on Context-Aware Guidelines
@@
Our work is motivated by the increasing availability of offline experiences that agents or humans naturally accumulate through their interactions with the environment. AutoGuide aims to leverage this offline data to improve the decision-making of an LLM agent by generating helpful context-aware guidelines. This section details how AutoGuide automatically constructs these guidelines and applies them to guide action generation at test time.
Figure 2: Context-aware guideline generation process based on a pair of contrastive trajectories $\bm{\tau}^{i}_{+}$ and $\bm{\tau}^{i}_{-}$. In this example, the two trajectories start deviating from each other at $t\!=\!0$. The context identification module generates a description of the context at $t\!=\!0$ given $\bm{\tau}^{i}_{:0}$, and the guideline extraction module generates the corresponding guideline for that context.
@@ 3.1 Problem Statement
Formally, AutoGuide is given offline data $\mathcal{D}_{\mathrm{train}}\!=\!(\bm{\tau}^{1},...,\bm{\tau}^{N})$ that consist of $N$ trajectories from training tasks. Each trajectory $\bm{\tau}\!=\!(x_{0},a_{0},r_{0},...,r_{T})$ is a sequence of observations, actions, and rewards following the partially observable Markov decision process [32]. The return of a trajectory is defined as the sum of rewards obtained throughout the trajectory: $R(\tau)\!=\!\textstyle\sum\nolimits_{t=0}^{T}{r_{t}}$. The objective of AutoGuide is to distill knowledge from offline experiences into a useful natural language format, such that the extracted information helps to maximize the expected return $\mathbb{E}_{\tau}[R(\tau)]$ during test time.
3.2 Extraction of Context-Aware Guidelines
@@
AutoGuide generates a set of context-aware guidelines by utilizing pairs of contrastive trajectories from offline data. Each context-aware guideline is expressed in concise natural language and follows a conditional structure, clearly describing the context in which the guideline is applicable. Intuitively, contrasting a pair of trajectories with different returns provides important information about when and which actions are effective or ineffective in maximizing expected returns. Building on this insight, we develop two modules for automatically extracting context-aware guidelines (see Figure 2):
Context identification module. This module is responsible for abstracting the given partial trajectory into its context, a concise natural language description of the agent’s state. More specifically, for a timestep $t$ and the corresponding trajectory $\bm{\tau}^{i}_{:t}\!:=\!(x_{0},a_{0},...,x_{t})$, we prompt LLMs to clearly describe the agent’s status:
$\begin{split}\textsc{context}\!\leftarrow\!\mathcal{M}_{\mathrm{context}}(\bm{\tau}^{i}_{:t}),\end{split}$ (1)
Our prompt templates for the context identification module are shown in Section C.1.
Algorithm 1 Extracting context-aware guidelines from offline data
@@
Input: Offline data $\mathcal{D}_{\mathrm{train}}$, context identification module $\mathcal{M}_{\mathrm{context}}$, guideline extraction module $\mathcal{M}_{\mathrm{guideline}}$
Initialize context-aware guideline dictionary $\mathcal{G}$
for Each pair $\bm{\tau}^{i}_{+},\bm{\tau}^{i}_{-}\in\mathcal{D}_{\mathrm{train}}$  do
   # Identify the context from a trajectory
   Find the deviating timestep $t$ from $\bm{\tau}^{i}_{+}$ and $\bm{\tau}^{i}_{-}$
   $\textsc{context}\leftarrow\mathcal{M}_{\mathrm{context}}(\bm{\tau}^{i}_{:t})$
   # Check if the current context matches any
    existing contexts
   if $\textsc{context}\notin\mathcal{G}$ then
      $\mathcal{G}[\textsc{context}]=\{\}$
@@
   # Generate the context-aware guideline
   $\textsc{guideline}\!\!\leftarrow\!\!\mathcal{M}_{\mathrm{guideline}}(\bm{\tau}^{i}_{+},\bm{\tau}^{i}_{-},\textsc{context})$
   $\mathcal{G}[\textsc{context}]\!\!\leftarrow\!\!\mathcal{G}[\textsc{context}]\cup\{\textsc{guideline}\}$
@@
Return Context-aware guideline dictionary $\mathcal{G}$
Algorithm 2 Applying context-aware guidelines at test time
Input: Context-aware guideline dictionary $\mathcal{G}$, context identification module $\mathcal{M}_{\mathrm{context}}$, guideline selection module $\mathcal{M}_{\mathrm{select}}$, LLM agent policy $\pi$
@@
   # Identify the current context from a trajectory
   $\textsc{context}\leftarrow\mathcal{M}_{\mathrm{context}}(\bm{\tau})$
   # If the current context matches any existing
    ones, perform top-$k$ guideline selection
   if $\textsc{context}\in\mathcal{G}$ then
      $\textsc{guidelines}\leftarrow\mathcal{M}_{\mathrm{select}}(\textsc{context},\bm{\tau};\mathcal{G},k)$
@@
      $\textsc{guidelines}\leftarrow\varnothing$
@@
   $a_{t}\sim\pi(\bm{\tau},\textsc{context},\textsc{guidelines})$
@@
   Update trajectory $\bm{\tau}\leftarrow\bm{\tau}\cup\{\textsc{context},a_{t},x_{t+1}\}$
@@
Guideline extraction module. This module aims to generate a desired guideline corresponding to a specific context. Let $\bm{\tau}^{i}_{+}$ and $\bm{\tau}^{i}_{-}$ represent a contrasting pair of trajectories for the same task $i$ in offline data $\mathcal{D}_{\mathrm{train}}$, where $R(\bm{\tau}^{i}_{+})\!>\!R(\bm{\tau}^{i}_{-})$. We want to contrast the pair of trajectories to find desired behaviors at an important timestep. To do this, we compare these two trajectories to find the deviation timestep $t$ at which they begin to diverge due to different actions. Then we apply the context identification module to summarize the context for the shared part of the trajectory $\bm{\tau}^{i}_{:t}$. Eventually, we extract a useful natural language guideline by examining the paired contrastive trajectories $\bm{\tau}^{i}_{+}$ and $\bm{\tau}^{i}_{-}$ with respect to the context:
$\begin{split}\text{{guideline}}\!\leftarrow\!\mathcal{M}_{\mathrm{guideline}}(\bm{\tau}^{i}_{+},\bm{\tau}^{i}_{-},\textsc{context}),\end{split}$ (2)
where we refer to Section C.2 for our prompt template. As an example, the paired trajectories in Figure 2 deviate from timestep $t=0$, for which the context is summarized as On the Reddit main page. This module then generates the following context-aware guideline: When on the Reddit main page, if you want to navigate to a specific forum, you should click on “link ‘Forums”’ located right above “link ‘Wiki”’.
@@
Construction of context-aware guidelines. We collect context-aware guidelines $\mathcal{G}$ by iterating through available pairs in the paired offline data and organize the guidelines in a dictionary format, using the context as the key and the corresponding guidelines as the value (see Algorithm 1). In particular, we observe that the context identification module occasionally produces contexts that describe the same situation but are expressed slightly differently. To minimize redundancy, we employ an LLM to determine if the current context corresponds to any previously identified context. If a match is found, we reuse the existing context; otherwise, we introduce a new context into our dictionary $\mathcal{G}$. The specific prompt template for this context-matching procedure is in Section C.3.
3.3 Applying Context-Aware Guidelines at Test Time